OpenStudy (anonymous):

can somebody help me graph the following equations and inequalities on a coordinate plane, shading the solution set where necessary. y= 4x-2

pooja195 (pooja195):

https://www.desmos.com/calculator

pooja195 (pooja195):

graph it using that and then test out your coordinates (0,0) is the easiest to do :)

OpenStudy (anonymous):

could you show me how I would work this out pooja?

pooja195 (pooja195):

sure how do you want me to graph it though?

OpenStudy (anonymous):

you could draw it if you wanted? I just would like to know how I would work the amswer out?

pooja195 (pooja195):

ok lets do this when graphing you always want to start with the Y axis so like this hold on...

pooja195 (pooja195):



pooja195 (pooja195):

and then you want to go up 4 over 1

OpenStudy (anonymous):

you got -2 because we had the y already and up to over one because we had the slope already?

pooja195 (pooja195):

-2 is on the Y axis it is not the slope

OpenStudy (anonymous):

I mean in the euation we have -2 and the slope is 4 because we had that already?

pooja195 (pooja195):

right but you have to graph it right?

OpenStudy (anonymous):

yes but I need to show my calculations as well

pooja195 (pooja195):

im so confused =/ @Directrix can you please help?

OpenStudy (anonymous):

i'm

OpenStudy (anonymous):

sorry to confuse you:[

Directrix (directrix):

I looked at this question before and I think that @Danny.Mcintire7 did not write the complete question as it appears in the text. >> help me graph the following equations and inequalities on a coordinate plane, shading the solution set where necessary. y= 4x-2 That is just one equations. Where are the other equations and the inequalities. So, @pooja195 the first thing we need to do is to find out what the complete question is.

OpenStudy (anonymous):

graph the following equations and inequalities on a coordinate plane, shading the solution where necessary. y=4x-2

OpenStudy (anonymous):

graph the following equation and inequalitie on a coordinate plane, shading the solution where necessary. y=4x-2 that is what im asked to do lol

pooja195 (pooja195):

@Directrix now im confused i dont know how to help this is all the info thats given =/

OpenStudy (anonymous):

so how would I graph the equation y=4x -2 on a graph and shade the solution where necessary?

Directrix (directrix):

What I get out of the question is to graph this: y=4x-2 That is the equation of a line with slope 4 and y-intercept -2. Two points determine a line One point on the line is (0, -2) because if x = 0, then y = 4*0 -2 = -2. A second point on the line is (1, 2) because if x = 1, then y = 4*1 - 2 = 2

Directrix (directrix):